# Quantum Field Theory on Multifractal Spacetime: Varying Dimension and Ultraviolet Completeness

### v1

@@ 3.2 Translations, rotations, and the vacuum
Let us first focus on the spatial translations which, as anticipated, play a central role in the proposed framework.
Similarly to the standard case, the 3-momentum operator is given by  [^4] The vacuum-to-vacuum expectation value of $P_{i}$ is zero in the standard QFT, but a priori not the one of $P_{0}$. Often, this is artificially made zero by Normal Product. One reason is that the corresponding integral rapidly diverges. This is in contrast with the QFT in the multifractal spacetime, where the finiteness does not require artifacts in the regularization.
@@
Since rotations are written as volume integrals of the type in equation (31), where one takes the textbook expressions and replaces $\phi(x)\to\phi(x,M)$ and then $a(\vec{p})\to a(\vec{p},M)$, the analog conclusion of equation (33) also holds for rotations. Thus the theory, due to the deformation of the generators of translation or rotations caused by $r$, shows a breaking at the quantum level of the Poincaré invariance.
@@ 4 Consistency of the theory
In this section, our proposal confronts two milestones of QFT: the Kallen-Lehmann representation of the propagator; the Gell-Mann and Low theorem for the propagator.
While the nonperturbative Kallen-Lehmann representation bounds the behavior of the propagator, thus providing a consistency check for the theory, the Gell-Mann and Low formula, base for perturbation theory, is not a test for the theory but is an improved result within nonstandard QFT including the scale $M$.
Notice that both these well-known results are related to the Poncaré (non)invariance of the vacuum. We shall now go through the issue in detail.
Kallen-Lehmann spectral representation.
In standard QFT, the Kallen-Lehmann representation of the propagator reads,
@@
where $\rho$ is a positive definite function, known as spectral function.
The above equation implies that the propagator cannot decay faster than $p^{-2}$ for large momentum, in apparent contrast with (29). However, the Kallen-Lehmann representation does not exist in the fractal QFT and this invalidates the above constraint. The existence of the Kallen-Lehmann representation in (35) requires the Poncaré invariance of the vacuum to manipulate the expression $\langle\Omega|\phi(x)\phi(y)|\Omega\rangle$. In particular, one has in the fractal QFT,
@@
In the first step, we used the completeness relation for a complete set $|n\rangle$, eigenvalues of the four-momentum, in the last step, we implemented the non-translational invariance of the vacuum, via the vacuum momentum $p_{\Omega}$. However, this is infinite since one has a finite momentum density in (32) and (33), corresponding to an infinite momentum for an infinite volume. Therefore, the expression in (36) becomes ill-defined.
We thus conclude that the Kallen-Lehmann representation of the two-point function does not exist if the vacuum is not invariant under translations, and therefore it does not constrain the expression (29) to behave as $1/p^{2}$ for large $p$.
@@
[…]
Standard QFT assumes the existence of the unitary operator connecting the free and interactive fields through:
@@
Moreover, the free and interactive vacuua are related as (calling $t^{\prime}=0$ and $t=T$, as $T\rightarrow\infty$) [27],
@@
[…]
@@ 5 Possible Implications and Discussion
In this section, we explore some implications of fractal QFT. The most striking consequence is perturbative finiteness – all loops become convergent. Moreover, we argue that perturbative series behave better at large order $n$ due to the absence of UV renormalons. Recall that renormalons lack any semi-classical interpretation that drives the factorial ($n!$) divergence in renormalized perturbation theory [29].
@@
$\left\{\begin{array}[]{ll}\lambda(\mu)\simeq\lambda(\mu_{0})+\left(1+\lambda(\mu_{0})\beta_{1}\log(\mu^{2}/\mu_{0}^{2})\right)&\mu<M\\[10.0pt] \lambda(\mu)\simeq\lambda(M)\left(1-\lambda(M)\beta_{1}\frac{M^{4}}{4\mu^{4}}e^{-\frac{\mu^{4}}{M^{4}}}\right)&\mu\geq M\,,\\ \end{array}\right.$ (41)
@@
In summary, any loop constructed in the model becomes finite, and no Landau pole exists due to (41).
@@
The superscripts -- and $+$ denote the creation and destruction operator parts of the field. The standard result is modified only by the factors $r$ in (42). Given the form of the function $r$ in (30), the part in $p^{2}$ becomes trivial onshell, $p^{2}=m^{2}$, while the part in $p_{\mu}v^{\mu}$ produces an anisotropy. Some momentum directions may be enhanced, depending on the constant four-vector $v$. Of course, all these modifications are suppressed at energies much lower than $M$.
@@
We have discussed the perturbative finiteness of the theory above, now let us consider what to expect at the non-perturbative level. Renormalons [29, 30] can be interpreted as a bridge between perturbative and non-perturbative physics. The singularities in the Borel transform due to renormalons lie on the semi-positive axis and hamper any Borel-Laplace resummation. They are not related to any semi-classical expansion and are considered genuine non-perturbative objects related to renormalization [30]. In the literature, their presence represents a failure of the perturbative renormalization procedure since the resulting series cannot be resummed without ambiguities [29].
Returning to the $\phi(x)^{4}$ model, renormalons relate to factorially divergent series [^6] Independently of the instantons that notwithstanding leading to a $n!$ large order behavior of the perturbation theory series, their existence does not imply any inconsistency in the theory, as explained in [29]. obtained by evaluating Feynman diagrams with many insertions of the same sub-diagram (the fish-diagram). The origin of the factorial $n!$ (where $n$ is the perturbation theory order) stems from the logarithmic high-energy behavior of the fish-diagram (sub-diagram) inserted $n$ times. While we refer to [31] for a complete derivation of the $n!$ renormalon behavior (see [32] for an alternative approach), here we illustrate its origin with a simple example. Consider the integral:
@@
This expression mimics a so-called bubble diagram, i.e. a loop integral with $n$ nested occurrences of the 4-point correlator (fish diagram), which has logarithmic UV behavior in standard QFT.
@@
It’s worth commenting on the nature of asymptotic safety shown in Fig. 1. Wilson first argued for the necessity of a non-trivial UV fixed point for theoretical consistency [37]. In our case – QFT with diminishing dimension – the UV fixed point is reached only at infinity. Nevertheless, the couplings remain finite and possibly small. While the fixed point is attained at infinite energy, the running exponentially converges to a constant value. Additionally, as explained previously for the $\phi^{4}$ model, the absence of UV renormalons enhances perturbation theory’s robustness.
Last but not least, as illustrated in (42), the multifractal QFT predicts direction-dependent cross-sections. Phenomenologically, within the standard model, this would probably be the most striking “smoking gun” signal of the underlying space-time fractality. In parallel, the Lorentz violating aspects of the proposed scenario should be phenomenologically bounded, being this kind of studies an active research field [38]. A detailed phenomenological analysis is beyond the scope of this work.
@@ 6 Conclusions and outlook
We have shown how to quantize fields in a spacetime with multifractal properties and how this automatically implies dimensional reduction, ensuring consistency with established field theory constraints. Notably, not only is quantization compatible with a varying dimension, but such variation significantly enhances QFT’s robustness, enabling rigorous $S$-matrix construction in the interaction picture (perturbation theory). This results in a finite theory without loop divergences and improves the perturbative series’ behavior. The significance of our contribution lies in demonstrating that non-differentiable spacetime structures are not only compatible with quantum field theory but actively improve its mathematical foundation. All this is achieved while retaining all known behaviors and predictions of standard QFT at low energies, yet predicting new behaviors at large energies, mainly through “asymptotic safety” and small anisotropies at high energy scatterings.
@@
In summary, while phenomenological studies in the literature investigate potential effects of varying dimensions [46], we believe our work provides the theoretical foundation for such scenarios. A dedicated study will be needed to analyze further consequences for particle physics and cosmology.
@@ Acknowledgement
AM thanks Emanuela Pichelli for discussions during this work on the appearance of fractals in the physical world.
@@ Appendix A S-matrix and Poincaré invariance of the vacuum
This appendix highlights some known problems in building a consistent $S-$matrix in interaction picture in QFT.
@@
In standard QFT, namely, in the absence of the fractality and reduction scale, the unitary operator $U(t_{0},t)$ exists only in the trivial case, with no interaction, for which such an operator is just the identity. This result is known as the Haag theorem, heavily relying on the translational invariance of the vacuum. It may be helpful to recall how this invariance leads to the no-go of the theorem, following [35] – see [47] for a rigorous proof within axiomatic QFT.
For shortness, we denote $U(t_{0},t):=U$ and call $T_{0}$ and $T$ the translational operator in the free and interactive Fock spaces. The translational invariance of the vacuua reads,

### v2

@@ 3.2 Translations, rotations, and the vacuum
Let us first focus on the spatial translations, which, as anticipated, play a central role in the proposed framework.
Similarly to the standard case, the 3-momentum operator is given by  [^5] The vacuum-to-vacuum expectation value of $P_{i}$ is zero in the standard QFT, but a priori not the one of $P_{0}$. Often, this is artificially set to zero by Normal Product. One reason is that the corresponding integral rapidly diverges. This is in contrast with the QFT in the multifractal spacetime, where the finiteness does not require artifacts in the regularization.
@@
Since rotations are written as volume integrals of the type in equation (36), where one takes the textbook expressions and replaces $\phi(x)\to\phi(x,M)$ and then $a(\vec{p})\to a(\vec{p},M)$, the analog conclusion of equation (38) also holds for rotations. Thus, the theory, due to the deformation of the generators of translation or rotations caused by $r$, shows a breaking at the quantum level of the Poincaré invariance.
The linear term $p\cdot v$ in the exponent of (35) represents a crucial departure from the form factors typically employed in nonlocal quantum gravity [30, 29, 31]. Standard nonlocal theories utilize form factors of the type $\exp(-p^{2}/M^{2})$, which depend only on $p^{2}$ and therefore preserve the Poincaré invariance of the vacuum. In contrast, our form factor explicitly breaks translational invariance through the (p/Mv)(p/M-v) structure, where $v$ is a constant four-vector that selects a preferred frame. This distinction is not merely technical but conceptually fundamental: while both approaches introduce exponential UV suppression, only the breaking of vacuum translational invariance—encoded in the $v$-dependence—circumvents Haag’s theorem (38), (39) and enables consistent construction of the interaction picture, as will be discussed in Sec.4. Nonetheless, standard nonlocal theories achieve UV finiteness and improved renormalization properties but remain subject to Haag’s constraints; our framework sacrifices exact Poincaré invariance of the vacuum to resolve the more fundamental obstruction to perturbative QFT posed by Haag’s theorem. The physical interpretation is that the multifractal structure of spacetime itself singles out a preferred direction encoded in $v$, with observable consequences in high-energy scattering cross-sections that vanish in the low-energy limit $p\ll M$.
@@ 4 Consistency of the theory
In this section, our proposal confronts two milestones of QFT: the Källén-Lehmann representation of the propagator; the Gell-Mann and Low theorem for the propagator – once an interaction, say $\frac{\lambda}{4!}\,\phi^{4}$, is turned on.
While the nonperturbative Källén-Lehmann representation bounds the behavior of the propagator, thus providing a consistency check for the theory, the Gell-Mann and Low formula, base for perturbation theory, is not a test for the theory but is an improved result within nonstandard QFT, including the scale $M$.
Notice that both these well-known results are related to the Poincaré (non)invariance of the vacuum. We shall now go through the issue in detail.
Källén-Lehmann spectral representation.
Recent work has shown that nonlocal theories with entire form factors and a Poincaré-invariant vacuum admit a generalized Källén–Lehmann representation with a positive spectral density and the correct local limit [55]. Our scenario is complementary: because the vacuum breaks translational invariance, a key step in the standard derivation fails. The usual KL bound does not apply to the discussed scenario.
In standard QFT, the Källén-Lehmann representation of the propagator reads,
@@
where $\rho$ is a positive definite function, known as the spectral function.
This equation implies that the propagator cannot decay faster than $p^{-2}$ for large momentum, in apparent contrast with (34). However, we argue that the Källén-Lehmann representation does not hold in the multifractal QFT, thereby invalidating the constraint. The key point is that deriving the Källén-Lehmann representation in (40) requires the Poincaré invariance of the vacuum to manipulate the expression $\langle\Omega|\phi(x)\phi(y)|\Omega\rangle$.
In particular, one has in the fractal QFT,
@@
In the first step, we used the completeness relation over a basis of four-momentum eigenstates $|n\rangle$; in the second, we incorporated the non-translational invariance of the vacuum via the vacuum momentum $p_{\Omega}$. The crucial point is that $p_{\Omega}$ is infinite since the vacuum carries a finite momentum density (see (37) and (38)). Therefore, the right-hand side of the expression in (41) becomes ill-defined, implying that the Källén-Lehmann representation of the two-point function breaks down in multifractal QFT. Consequently, the standard constraint that the propagator cannot decay faster than $1/p^{2}$ at large $p$ no longer applies.
@@
$U(t,t^{\prime})=\exp\left[-i\int_{t^{\prime}}^{t}H_{I}(\tau)d\tau\right]$ (42)
Standard QFT assumes the existence of the unitary operator connecting the free and interacting fields through:
@@
Moreover, the free and interactive vacua are related as (calling $t^{\prime}=0$ and $t=T$, as $T\rightarrow\infty$) [56],
@@
$\langle\Omega|\mathcal{T}\phi(x)\phi(y)|\Omega\rangle=\frac{\langle 0|\mathcal{T}\phi_{0}(x)\phi_{0}(y)\exp[-i\int_{-\infty}^{\infty}H_{I}(t)dt]|0\rangle}{\langle 0|\mathcal{T}\exp[-i\int_{-\infty}^{\infty}H_{I}(t)dt]|0\rangle}$ (45)
@@ 5 Possible Implications and Discussion
In this section, we explore some implications of fractal QFT. The most striking consequence of the adopted form factor is perturbative finiteness. All loops become convergent, and the running coupling ceases to grow logarithmically at high energies, eliminating UV renormalons. Recall that renormalons lack any semi-classical interpretation that drives the factorial ($n!$) divergence in renormalized perturbation theory [58].
@@
[…]
@@
In summary, any loop constructed in the model becomes finite, and no Landau pole exists due to (46). This points to super-renormalizability or even finiteness, properties extensively studied in nonlocal quantum gravity [30, 52, 33, 24, 36]. The exponential form factor suppresses UV divergences at all loop orders.
Anisotropies in high-energy scatterings.
@@
The superscripts - and ++ denote the creation and destruction operator parts of the field. The standard result is modified only by the factors $r$ in (47). Given the form of the function $r$ in (35), the part in $p^{2}$ becomes trivial on-shell, $p^{2}=m^{2}$, while the part in $p_{\mu}v^{\mu}$ produces an anisotropy. Some momentum directions may be enhanced, depending on the constant four-vector $v$. Of course, all these modifications are suppressed at energies much lower than $M$.
@@
We have discussed the perturbative finiteness of the theory above; now, let us consider what to expect at the non-perturbative level. Renormalons [58, 59] can be interpreted as a bridge between perturbative and non-perturbative physics. The singularities in the Borel transform due to renormalons lie on the semi-positive axis and hamper any Borel-Laplace resummation. They are not related to any semi-classical expansion and are considered genuine non-perturbative objects related to renormalization [59]. In the literature, their presence represents a failure of the perturbative renormalization procedure since the resulting series cannot be resummed without ambiguities [58].
Returning to the $\phi(x)^{4}$ model, renormalons relate to factorially divergent series [^7] Independently of the instantons that, notwithstanding leading to a $n!$ large order behavior of the perturbation theory series, their existence does not imply any inconsistency in the theory, as explained in [58]. obtained by evaluating Feynman diagrams with many insertions of the same sub-diagram (the fish-diagram). The origin of the factorial $n!$ (where $n$ is the perturbation theory order) stems from the logarithmic high-energy behavior of the fish-diagram (sub-diagram) inserted $n$ times. While we refer to [60] for a complete derivation of the $n!$ renormalon behavior (see [61] for an alternative approach), here we illustrate its origin with a simple example. Consider the integral:
@@
This expression mimics a so-called bubble diagram, i.e., a loop integral with $n$ nested occurrences of the 4-point correlator (fish diagram), which has logarithmic UV behavior in standard QFT.
@@
It’s worth commenting on the nature of safety shown in Fig. 1. Wilson first argued for the necessity of a non-trivial UV fixed point for theoretical consistency [66]. In our case – QFT with diminishing dimension – the UV fixed point is reached only at infinity. Nevertheless, the couplings remain finite and possibly small. While the fixed point is attained at infinite energy, the running exponentially converges to a constant value.
The realization of asymptotic safety discussed in this work differs fundamentally from Weinberg’s asymptotic safety program [67, 68, 69, 70, 71] and its implementations in functional renormalization group (FRG) approaches to quantum gravity [72, 73, 74]. In standard asymptotic safety, UV finiteness arises from a non-Gaussian fixed point of the gravitational RG flow where dimensionless couplings become scale-independent, with dimensional reduction emerging as a consequence of the fixed-point structure [75, 76]. The theory remains perturbatively non-renormalizable but is well-defined non-perturbatively through the fixed point, preserving Poincaré invariance throughout. In contrast, our framework achieves asymptotic safety through the exponential form factor suppression $e^{-p^{2}/M^{2}}$ in (35), which causes the running coupling to asymptotically approach a constant (46) without reaching a genuine RG fixed point—the coupling continues to vary, albeit exponentially slowly, rather than becoming exactly scale-independent. Moreover, our mechanism operates entirely within an improved perturbation theory (due to Haag’s no-go evasion and a coupling that can remain small during the running). While both approaches point to dimensional reduction at high energies, the physical mechanisms differ. The standard asymptotic safety relies on strong quantum gravity fluctuations driving the theory to a fixed point, whereas the approach presented here accomplishes the UV completion through the form factor, effectively parameterizing the flow of the spacetime dimension.
As illustrated in (47), the multifractal QFT predicts direction-dependent cross-sections. Phenomenologically, within the standard model, this would probably be the most striking “smoking gun” signal of the underlying space-time fractality. This breaking of the vacuum’s translational and rotational invariance, encoded in the vector $v$, implies that Lorentz invariance is also broken at a fundamental level. This is a crucial feature of our model with significant phenomenological implications. While a detailed analysis of experimental bounds is beyond the scope of this work [77], it is essential to address a key theoretical consistency check for such theories.
A critical challenge for Lorentz-violating theories is the potential for radiative corrections to amplify symmetry-breaking effects to phenomenologically unacceptable levels [78, 79]. Crucially, our framework is protected from this instability due to the term $e^{-p^{2}/M^{2}}$ in the form factor of (35). The amplification mechanism is a direct consequence of having a form factor of the form $F(p\cdot v)$, with $F(p\cdot v)=1$ for $p\rightarrow 0$ and $F(p\cdot v)=0$ for $p\rightarrow\infty$. In our model, the form factor in (35) ensures renormabizability is preserved, namely unsupressed contributions proportional to positive powers of the scale $M$ can be reabsorbed into a redefinition of the couplings present in the tree level Lagrangian. Thus any Lorentz-violating contribution must be suppressed by negative powers of the large scale $M$. This eliminates the source of any pathological enhancement and ensures decoupling. See Appendix B for a detailed and explicit discussion.
@@ 6 Conclusions and outlook
We have shown how to quantize fields in a spacetime with multifractal properties and how this automatically implies dimensional reduction, ensuring consistency with established field theory constraints. Notably, not only is quantization compatible with a varying dimension, but such variation significantly enhances QFT’s robustness, enabling rigorous $S$-matrix construction in the interaction picture (perturbation theory). This results in a finite theory without loop divergences and improves the perturbative series’ behavior. The significance of our contribution lies in demonstrating that non-differentiable spacetime structures are not only compatible with quantum field theory but also actively improve its mathematical foundation. All this is achieved while retaining all known behaviors and predictions of standard QFT at low energies, yet predicting new behaviors at large energies, mainly through “asymptotic safety” and small anisotropies at high-energy scatterings.
@@
In summary, while phenomenological studies in the literature investigate potential effects of varying dimensions [86, 87, 88, 89], we believe our work provides the theoretical foundation for such scenarios. A dedicated study will be needed to analyze further consequences for particle physics and cosmology.
@@ Acknowledgement
AM thanks Emanuela Pichelli for discussions during this work on the appearance of fractals in the physical world. JCV give special thanks to Paolo Creminelli , Marco Serone, and Andrea Romanino for support. JCV also gives special thanks to SISSA International Relations Office, in particular to Francesca Gandolfo and Gabriella Perissini for help and guidance.
@@ Appendix A S-matrix and Poincaré invariance of the vacuum
This appendix highlights some known problems in building a consistent $S$-matrix in the interaction picture in QFT.
@@
In standard QFT, namely, in the absence of fractality and reduction scale, the unitary operator $U(t_{0},t)$ exists only in the trivial case, with no interaction, for which such an operator is just the identity. This result is known as the Haag theorem, heavily relying on the translational invariance of the vacuum. It may be helpful to recall how this invariance leads to the no-go of the theorem, following [64] – see [90] for a rigorous proof within axiomatic QFT.
For shortness, we denote $U(t_{0},t):=U$ and call $T_{0}$ and $T$ the translational operators in the free and interactive Fock spaces. The translational invariance of the vacua reads,
@@
Appendix B Quantum corrections do not enhance Lorentz violation
In this appendix, we argue that the framework presented in this paper avoids the pathologies pointed out in [78, 79] regarding loop-enhanced Lorentz asymmetries that rule out the phenomenological viability of some models.
Consider the one-loop correction to the self-energy $\Pi$ in the $\phi^{4}$ model, embedded in the setting of this work, namely with the propagator modified by the form factor in (35). The one-loop contribution reads (in Euclidean momentum),
$\Pi=\frac{\lambda}{2}\int\frac{d^{4}l_{E}}{(2\pi)^{4}}\frac{e^{-l_{E}^{2}/M^{2}-l\cdot v/M}}{l_{E}^{2}+m^{2}}\,.$ (63)
Since we are interested in the UV behavior, we evaluate the integral in the limit $m=0$. We write $l\cdot v=|l|\times|v|\cos(\theta)$ and rewrite the 4D Euclidean integration as
$\int_{0}^{\infty}\int_{0}^{\pi}\int_{0}^{\pi}\int_{0}^{2\pi}\frac{e^{-l_{E}^{2}/M^{2}-l\cdot v/M}}{l_{E}^{2}}l_{E}^{3}\sin(\theta)^{2}\sin(\varphi)dl_{E}d\theta d\varphi d\omega\,.$ (64)
The momentum integration leads to:
$\frac{1}{4}M^{2}\left(2-\sqrt{\pi}v\cos(\theta)e^{\frac{1}{4}v^{2}\cos^{2}(\theta)}\text{erfc}\left(\frac{1}{2}v\cos(\theta)\right)\right)\,.$ (65)
where erfc is the complementary error function.
Next, we proceed to the $\theta$-integration. Here comes the crucial point: upon integration over $\theta$, the term linear in $v\cdot\cos(\theta)$ vanishes due to odd parity. Therefore, only even power of $v$ can appear in the result.
More explicitly, after the trivial integration in $\varphi$ and $\omega$, the final result can be approximated as
$\Pi\simeq\frac{\lambda M^{2}\left(v^{4}+12v^{2}+96\right)}{3072\pi^{2}}+O\left(v^{5}\right)\,.$ (66)
Thus, we have $\Pi$, resulting in a large finite term (due to the presence of $M$) that can be reabsorbed into the standard counterterms, specifically the mass counterterm. Note that the form factor provides regularization, but renormalization is still at work, and then these terms do not lead to unsuppressed Lorentz violating terms.
To make contact with [78, 79], suppose we had a form factor as
$e^{-l\cdot v/M}\,$,
now without the $l^{2}/M^{2}$ suppression. In this case, the integral corresponding to the one in (63) would not automatically be finite. To see this, rewrite $l_{E}=l_{E}^{p}+l_{E}^{t}$, with $l_{E}^{p},l_{E}^{t}$ being the parallel and transverse components with respect to the vector $v$. As a result, only the integral over $l_{E}^{p}$ would be finite, while those over $l_{E}^{t}$ (3 dimensions) would require a cutoff since $v\cdot l_{E}^{t}=0$. This would then be a direction-dependent cutoff in 3 dimensions, similar to the one discussed in Sect. 1.3 of [79]. In this example, one would then have enhanced Lorentz violation via the Lorentz-asymmetric cutoff. Thus, the crucial role in our framework—unlike in [78, 79]—is played by the term $\exp(-p^{2}/M^{2})$, which makes the loop finite in any direction in momentum space.
The arguments above apply in general. Consider a generic expression for a loop calculation of the two-point Green function ($p^{2}\ll M^{2}$):
$\Delta\Pi(p,v,m=0,M)=$ $\int\prod_{\ell=1}^{m}d^{4}\,\ell_{l}\,\mathcal{I}(\ell_{l},p,v,m,M)=A\,M^{2}+$
$B\,p^{2}+C\,\left(\frac{p\cdot v}{M}\right)\,p^{2}+\mathcal{O}\left(\frac{p^{n+2}}{M^{n}}\right)\,.$ (67)
where $A$, $B$, $C$ are functions of $\frac{p^{2}}{M^{2}}$ and $\frac{v\cdot p}{M}$. Since the renormalization properties of the theory are preserved, the Lorentz-violating finite contributions due to $v$ are automatically suppressed by the large scale $M$, unlike the potentially large Lorentz-violating contributions pointed out in [78, 79].